If the sum of a number and 6 is doubled, the result is 7 less than the number. Find the number
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39620)
Answer by ikleyn(52804)

.
The literal translation of the condition to an equation is this


    2*(x + 6) = x - 7.


Simplify and solve for x:


    2x + 12 = x - 7

    2x - x = -7 - 12 


    x = -19.


Answer.  The number under the question is -19.


Check.  The sum of the number and 6 is  -19+6 = -13.

        When doubled,  -13 becomes -26.

        7 less than the number is  -19 - 7 = - 26.

        ! Correct !

Solved and checked.